Solve for x
4x+3/1-2x=3/8

Answer by rfer(16322)

2x=3/8-3/1
2x=3/8-24/8
2x=-21/8
x=(-21/8)/(2)
x=-21/8*8/16
x=-168/128
x=-1 5/16